# FT: Unlocking the Power of Code Instruction Tuning by Simply Merging Upcycled Mixture-of-Experts

### v1

@@ Abstract
We introduce FT, a simple yet powerful training scheme, by simply merging upcycled Mixture-of-Experts (MoE) to unleash the performance limit of instruction-tuned code Large Language Models (LLMs). While vanilla sparse upcycling fails to improve instruction tuning, FT introduces a shared expert mechanism with a novel routing weight normalization strategy into sparse upcycling, which significantly boosts instruction tuning. After fine-tuning the upcycled MoE model, FT introduces a learnable model merging mechanism to compile the upcycled MoE back to a dense model, achieving upcycled MoE-level performance with only dense-model compute. By applying FT to a 1.3B model, we create a new state-of-the-art tiny code LLM (<3B) with 67.1 and 64.6 pass@1 on HumanEval and HumanEval+ respectively. With the same data and model architecture, FT improves supervised fine-tuning (SFT) by 13% on HumanEval+, along with consistent improvements from 2% to 13% on MBPP+, MultiPL-E, and DS-1000, demonstrating its generalizability. FT is fully orthogonal to existing techniques such as Evol-Instruct and OSS-Instruct, opening a new dimension for improving code instruction tuning. Codes are available at https://github.com/ise-uiuc/xft.
@@ 1 Introduction
A typical instruction tuning flow involves two steps Zhang et al. (2023): (i) curating an instruction dataset of instruction-output pairs, where the instruction reflects human intents in natural language and the output includes explained code snippets that correspond to the intent; and (ii) supervised fine-tuning of pre-trained LLM on the instruction dataset. In the realm of code, multiple instruction-tuning methods have been proposed to curate high-quality instruction datasets. For example, Code Evol-Instruct Luo et al. (2023) uses ChatGPT to obtain complex synthetic code instructions with heuristic prompts, while OSS-Instruct Wei et al. (2023) prompts ChatGPT to generate new coding problems by drawing inspiration from open source code snippets. While existing work focuses on the data perspectives of instruction tuning, they all follow the standard SFT, leaving room for exploring advanced training schemes.
@@
We argue that prior works largely overlook the possibility of improving the code instruction tuning by advancing the training schemes. Figure 1 depicts the supervised fine-tuning (SFT), which directly starts with the pre-trained weights and architecture for fine-tuning. The model is dense here because all parameters are activated to compute the next token (assuming it is a decoder-only LLM). In contrast to fine-tuning a dense model, following the scaling laws Kaplan et al. (2020) (i.e., more parameters, better performance), sparse upcycling Komatsuzaki et al. (2023) is proposed to efficiently upgrade the model sizes by upcycling a dense LLM to a sparsely activated Mixture-of-Experts (MoE) model. An MoE model is efficient because the generation of the next token only involves a subset of parameters (i.e., experts) and thus is sparsely activated. For example, Mixtral-8x7B Jiang et al. (2024), compared to a dense 7B model, uses approximately parameters and computation, i.e., only 2 out of 8 experts are dynamically selected to compute the next token. However, there are two key limitations when using sparse upcycling in instruction tuning: (i) Slow scaling: Komatsuzaki et al. (2023) show that sparse upcycling improves the dense SFT marginally at the early phase, requiring orders of magnitude of extra compute to achieve decent improvement; and (ii) Inference cost: though MoE is more efficient than directly scaling the size of dense LLMs, MoE is still expensive, especially at inference, as it introduces significantly more parameters (i.e., memory) and, more importantly, computes during inference, compared to its dense counterparts.
@@
In this paper, we propose FT: by simply merging upcycled MoE models, we push the performance limit of instruction-tuned code LLMs. While vanilla sparse upcycling fails to improve instruction tuning efficiently Komatsuzaki et al. (2023), FT addresses this challenge by isolating one expert as the shared expert among all the other experts in each MoE layer, inspired by DeepSeekMoE Dai et al. (2024) and MoCLE Gou et al. (2024). FT also includes a novel routing weight normalization strategy to eliminate scale mismatch between the upcycled MoE layer with the shared expert and the original dense layer, which will otherwise lead to performance degradation Wu et al. (2022). After the upcycled MoE model finishes the SFT phase, motivated by Model Soups Wortsman et al. (2022), FT uses a learnable model merging mechanism to output a dense model by merging all the expert networks in the upcycled MoE, i.e., the final dense model is of the same model structure and size as the original pre-trained model, achieving similar performance without paying extra inference cost as the sparse upcycling. With only 1.3B parameters, FT achieves 67.1 pass@1 on HumanEval and 64.6 pass@1 on HumanEval+, which is the new state-of-the-art for tiny code LLMs (<3B). Compared with SFT, FT achieves 13% improvement on HumanEval+. Surprisingly, our model merging mechanism can preserve or even further boost the general performance of the upcycled MoE with around parameters! We conclude our contribution as follows:
@@
Technique: We present FT, a new training scheme for code instruction tuning. FT involves two steps: upcycling and merging. A pre-trained dense LLM is first upcycled into an MoE with the shared expert setting and then fine-tuned on the instruction dataset. To avoid the performance degradation caused by the scale mismatch issue, we propose a novel routing weight normalization strategy. In addition, we introduce the first learnable mechanism for merging the upcycled MoE into a dense model, eliminating additional inference overhead while preserving or even improving the MoE performance.
Results: With only 1.3B parameters, FT achieves 67.1 pass@1 on HumanEval and 64.6 pass@1 on HumanEval+, which is the new state-of-the-art for tiny code LLMs (<3B). Compared with normal supervised fine-tuning (SFT), FT achieves 13% improvement on HumanEval+! FT also achieves a consistent improvement from 2% to 13% on MBPP, MultiPL-E, and DS-1000 over SFT, demonstrating its generalization.
@@ 2.1 Mixture-of-Experts
where refers to the total number of experts, refers to the gate value for the -th expert, refers to the -th expert, refers to the hidden states of the -th token which is the input of the -th MoE layer, refers to the affinity score between the -th expert and the -th token, refers to a function computing largest scores over , and refers to the centroid of the -th expert in the -th MoE layer. By definition, each token will only be assigned to and computed in the top experts among all the experts.
Recently, many works have been proposed to scale model sizes with MoE architecture Lepikhin et al. (2020); Du et al. (2022); Fedus et al. (2022); Jiang et al. (2024); Xue et al. (2024). While most MoE models are trained from scratch, sparse upcycling Komatsuzaki et al. (2023) is proposed to initialize MoE models based on the pre-trained dense model, which can efficiently reduce the computational costs of training MoE models, compared with training MoE models from scratch. Specifically, sparse upcycling constructs a new MoE model by initializing each expert of each MoE layer as a copy of the original FFN layer in the dense model, while directly copying the remaining layers from the dense model to the new MoE model.
@@ 2.2 Instruction Tuning
Instruction tuning is designed to improve the instruction-following ability of LLMs by fine-tuning them on the instruction datasets in a supervised fashion Wei et al. (2022). The quality of the instruction dataset is significant for the effectiveness of instruction tuning and researchers have proposed multiple methods to improve data quality. For example, Self-Instruct Wang et al. (2023) synthesizes high-quality instruction data by prompting a foundation LLM with specially designed prompts. To improve Self-Instruct, Evol-Instruct Xu et al. (2023) improves the complexity and diversity of the instruction dataset by prompting ChatGPT with heuristic prompts. OSS-Instruct Wei et al. (2023) queries ChatGPT to generate instruction-output pairs by getting inspiration from real-world code snippets.
Recently, some parameter-efficient fine-tuning techniques have been proposed to use MoE for better instruction tuning. For example, LoRAMoE Dou et al. (2023) and MoCLE Gou et al. (2024) propose MoE-like modules that are constructed with Low-Rank Adaptations (LoRA) to improve instruction tuning, while PESC Wu et al. (2024) proposes to integrate adapters into MoE that are upcycled from dense models. Different from these works, FT focuses on full fine-tuning, which generally performs better than parameter-efficient fine-tuning Chen et al. (2022).
@@ 2.3 Weight Averaging
Weight averaging is a commonly used technique to improve the performance of deep learning models. For example, Model Soups Wortsman et al. (2022) averages the weights of multiple models that are initialized from the same pre-trained model but finetuned with different hyperparameter configurations to improve the accuracy and robustness of the model. However, only a few works have been proposed to merge expert networks of an MoE layer to a normal FFN layer using weight averaging. For example, OneS Xue et al. (2022) proposes several simple weight averaging methods to merge expert networks of a BERT-based MoE model. Closely related to our work, Experts Weights Averaging (EWA) Huang et al. (2023) proposes to convert an MoE model to a dense model with two steps: (i) During MoE training, EWA conducts weighted averaging of all the expert weights after each weight update of MoE, which is based on a manually-crafted hyperparameter ; (ii) After training, EWA converts each MoE layer into an FFN layer by uniformly averaging the experts.
Different from all the aforementioned existing works, FT is the first work proposing a learnable mechanism to merge expert networks in the upcycled MoE model. Furthermore, while the training scheme of EWA is deeply coupled to a specific MoE architecture, FT can be easily adapted to different MoE architectures by only adjusting the final merging process. In addition, unlike EWA, FT does not introduce any hyperparameters into the training of the large MoE models, significantly reducing the computational resources for hyperparameter searching. Our empirical results in Section 4 also showcase the clear advantage of FT.
@@ 3.1 Upcycling
Inspired by sparse upcycling Komatsuzaki et al. (2023), we convert the pre-trained dense LLM to a new MoE by initializing each expert of each MoE layer as a copy of the original FFN layer in the dense model, while directly copying the remaining layers from the dense model to the new MoE model. However, the performance gain brought by sparse upcycling is negligible with a very limited extra training budget Komatsuzaki et al. (2023) – which is exactly the situation we are facing during instruction tuning. Intuitively, it is because each expert in the upcycled MoE model is trained on fewer instruction data than the original dense model does because traditional routers used in sparse upcycling will assign different tokens to different experts and thus reduce the amount of data each expert is trained on Gou et al. (2024). Consequently, inspired by DeepSeekMoE Dai et al. (2024) and MoCLE Gou et al. (2024), FT introduces the shared expert setting into sparse upcycling to tackle this challenge. We further propose a novel routing weight normalization strategy for FT to avoid the potential performance degradation caused by the scale mismatch problem Wu et al. (2022).
@@ 3.1.1 Shared Expert for Upcycling
During upcycling, we isolate one shared expert among all the other normal experts in each MoE layer, where the shared expert will be deterministically assigned to handle all the tokens while other normal experts are assigned by the router. By doing so, the upcycled MoE model can achieve a clear performance boost in instruction tuning, where the shared expert can learn general knowledge across the whole instruction dataset while other normal experts learn specific knowledge among different instructions assigned by the router. Formally, the output hidden state of the -th MoE layer when processing the -th token can be expressed as:
where refers to the output hidden state of the -th MoE layer when processing the -th token, refers to the total number of experts, refers to the gate value for the -th expert, refers to the -th expert, refers to the output hidden state of the -th attention layer when processing the -th token and also the input of the -th MoE layer, refers to the affinity score between the -th expert and the -th token, refers to the maximum affinity score among all the experts besides the shared expert, refers to a function computing largest scores over , refers to a set of largest affinity scores among all the experts besides the shared expert, and refers to the centroid of the -th expert in the -th MoE layer.
@@
Routing Weight Normalization. Although the shared expert setting is also used in recent works Dai et al. (2024); Gou et al. (2024), we cannot directly follow their routing strategy because they cannot handle a scale mismatch problem that is unique for sparse upcycling. The scale mismatch problem is that differences between the scale of the output of the upcycled MoE layer and the original FFN layer can cause performance degradation Wu et al. (2022). To handle this problem, we need to make sure the sum of equals 1, so that the output of the MoE layer matches that of the FFN layer in scale. To do so, we normalize the affinity scores of top normal experts with Softmax and scale their sum to to make sure that the sum of the of top experts, including one shared expert and normal experts, equals 1.
@@ 3.2 Merging
We propose a learnable model merging method to convert the large MoE model, namely MoE, back to a dense model FT. By doing so, we expect FT to keep the boosted performance gained during upcycling while keeping its model size the same as the original dense model to avoid any additional inference overhead. Inspired by Model Soups Wortsman et al. (2022), we choose to merge MoE by learning the mixing coefficients that can be used to average the parameters of all experts in each MoE layer to obtain a normal FFN layer, while directly copying other remaining layers.
Formally speaking, given the weights of experts at the -th layer , the process of merging each MoE layer to an FFN layer can be stated as below:
where denotes the merged parameter of all experts and denotes the learnable mixing coefficient of expert . We consider a neural network with input and parameters . For loss and instruction dataset , such mixing coefficients of all the layers can be learned via:
where refers to all the remaining layers of MoE other than MoE layers and is parameterized as the output of a softmax, so that each is positive and .
While the learning process defined in Eq. (4) is the most intuitive way of learning , our experiment in Section 5.2 shows that, due to the shared expert setting, it tends to simply increase the mixing coefficient of the shared expert at each layer as much as possible to decrease the loss. It is not helpful because, although the shared expert has learned general knowledge across the whole instruction dataset and needs a relatively large mixing coefficient, we still need to keep the scale of the mixing coefficient of other normal experts at a certain level also to keep some specific knowledge learned by other normal experts in the merged parameter .
To solve this issue, we introduce a shared expert rate to fix the mixing coefficient of the shared expert and learn the mixing coefficients of the remaining normal experts which sums to in each layer. By doing so, we can easily control the scale of the mixing coefficient of the shared expert, while still being able to learn the optimal layer-wise mixing coefficients of other normal experts. Let’s say is the shared expert of the -th layer, then Eq. (3) and Eq. (4) can be reformulated as below:
In practice, we uniformly initialize the mixing coefficients of all the normal experts as , which is then trained on the same instruction dataset as upcycling.

### v2

@@ Abstract
We introduce FT, a simple yet powerful training scheme, by simply merging upcycled Mixture-of-Experts (MoE) to unleash the performance limit of instruction-tuned code Large Language Models (LLMs). While vanilla sparse upcycling fails to improve instruction tuning, FT introduces a shared expert mechanism with a novel routing weight normalization strategy into sparse upcycling, which significantly boosts instruction tuning. After fine-tuning the upcycled MoE model, FT introduces a learnable model merging mechanism to compile the upcycled MoE model back to a dense model, achieving upcycled MoE-level performance with only dense-model compute. By applying FT to a 1.3B model, we create a new state-of-the-art tiny code LLM (<3B) with 67.1 and 64.6 pass@1 on HumanEval and HumanEval+ respectively. With the same data and model architecture, FT improves supervised fine-tuning (SFT) by 13% on HumanEval+, along with consistent improvements from 2% to 13% on MBPP+, MultiPL-E, and DS-1000, demonstrating its generalizability. FT is fully orthogonal to existing techniques such as Evol-Instruct and OSS-Instruct, opening a new dimension for improving code instruction tuning. Codes are available at https://github.com/ise-uiuc/xft.
@@ 1 Introduction
A typical instruction tuning flow involves two steps Zhang et al. (2023): (i) curating an instruction dataset of instruction-output pairs, where the instruction reflects human intents in natural language and the output includes target code snippets that correspond to the intent; and (ii) supervised fine-tuning of pre-trained LLM on the instruction dataset. In the realm of code instruction tuning, most recent works have been focusing on curating high-quality instruction datasets. For example, Code Evol-Instruct Luo et al. (2023) uses ChatGPT to obtain complex synthetic code instructions with heuristic prompts, while OSS-Instruct Wei et al. (2023) prompts ChatGPT to generate new coding problems by drawing inspiration from open source code snippets. Since existing works focus on the data perspectives of instruction tuning, they all follow the standard SFT, leaving room for exploring advanced training schemes.
@@
We argue that prior works largely overlook the possibility of improving code instruction tuning by advancing existing training schemes. Figure 1 depicts supervised fine-tuning (SFT), which directly uses the pre-trained weights and architecture for fine-tuning. The model is dense here because all parameters are activated to predict the next token (assuming it is a decoder-only LLM). In contrast to fine-tuning a dense model, following the scaling laws Kaplan et al. (2020) (i.e., more parameters, better performance), sparse upcycling Komatsuzaki et al. (2023) is proposed to efficiently upgrade the model size by upcycling a dense LLM to a sparsely activated Mixture-of-Experts (MoE) model. An MoE model is efficient because its prediction of the next token only invokes a subset of parameters (i.e., experts) and thus is sparsely activated. For example, Mixtral-8x7B Jiang et al. (2024), compared to a dense 7B model, uses approximately parameters and computation, i.e., only 2 out of 8 experts are dynamically selected to compute the next token. However, there are two key limitations when using sparse upcycling in instruction tuning: (i) Slow scaling: it is reported that sparse upcycling improves the performance of dense models marginally with limited training steps, requiring orders of magnitude of extra compute to achieve decent improvement Komatsuzaki et al. (2023); (ii) Inference cost: although MoE is more efficient than directly scaling up the size of dense LLMs, MoE is still expensive, especially at inference, as it introduces significantly more parameters (i.e., memory) and computes during inference, compared to its dense counterparts.
@@
In this paper, we propose FT: by simply merging upcycled MoE models, we push the performance limit of instruction-tuned code LLMs. While vanilla sparse upcycling fails to improve instruction tuning efficiently Komatsuzaki et al. (2023), FT addresses this challenge by isolating one expert as the shared expert among all the other experts in each MoE layer, inspired by DeepSeekMoE Dai et al. (2024) and MoCLE Gou et al. (2024). FT also proposes a novel routing weight normalization strategy to eliminate scale mismatch between the upcycled MoE layer with the shared expert and the original dense layer, which will otherwise lead to performance degradation Wu et al. (2022). After the upcycled MoE model finishes its SFT phase, motivated by Model Soups Wortsman et al. (2022), FT uses a learnable model merging mechanism to output a dense model by merging all the expert networks in the upcycled MoE, i.e., the final dense model is of the same model structure and size as the original pre-trained model, achieving similar performance without paying extra inference cost as the sparse upcycling. With only 1.3B parameters, FT achieves 67.1 pass@1 on HumanEval and 64.6 pass@1 on HumanEval+, which is the new state-of-the-art for tiny code LLMs (<3B). Compared with SFT, FT achieves 13% improvement on HumanEval+. Surprisingly, our learnable merging mechanism can preserve or even further boost the performance of the upcycled MoE with only around parameters! We conclude our contribution as follows:
@@
Technique: We present FT, a new training scheme for code instruction tuning. FT involves two steps: upcycling and merging. A pre-trained dense LLM is first upcycled into an MoE with the shared expert setting and then fine-tuned on the instruction dataset. We propose a novel routing weight normalization strategy to avoid the performance degradation caused by the scale mismatch problem. In addition, we introduce the first learnable mechanism for merging the upcycled MoE into a dense model, eliminating additional inference overhead while preserving or even improving the upcycled MoE performance.
Results: With only 1.3B parameters, FT achieves 67.1 pass@1 on HumanEval and 64.6 pass@1 on HumanEval+, which is the new state-of-the-art for tiny code LLMs (<3B). Compared with normal SFT, FT achieves a significant 13% improvement on HumanEval+! FT also achieves consistent improvements from 2% to 13% on MBPP, MultiPL-E, and DS-1000 over SFT, demonstrating its generalizability.
@@ 2.1 Mixture-of-Experts
where refers to the gate value for the -th expert given the -th token, refers to the -th expert, refers to the hidden states of the -th token which is the input of the -th MoE layer, refers to the affinity score between the -th expert and the -th token, refers to a function extracting highest scores out of , and refers to the centroid of the -th expert in the -th MoE layer. By definition, each token will only be computed in the top experts among all the experts and such sparsity assures the efficiency of MoE.
Recently, many works have been proposed to scale model sizes with MoE architecture Lepikhin et al. (2020); Du et al. (2022); Fedus et al. (2022); Jiang et al. (2024); Xue et al. (2024). While most MoE models are trained from scratch, sparse upcycling Komatsuzaki et al. (2023) is proposed to initialize MoE models based on pre-trained dense models, which can efficiently reduce the computational costs of training MoE models, compared with training MoE models from scratch. Specifically, sparse upcycling constructs a new MoE model by initializing each expert of each MoE layer as a copy of the original FFN layer in the dense model, while directly copying the remaining layers from the dense model to the new MoE model.
@@ 2.2 Instruction Tuning
Instruction tuning is designed to improve the instruction-following ability of LLMs by fine-tuning them on the instruction datasets in a supervised fashion Wei et al. (2022). The quality of the instruction dataset is significant for the effectiveness of instruction tuning and researchers have proposed multiple methods to improve data quality. For example, Self-Instruct Wang et al. (2023) synthesizes high-quality instruction data by prompting a foundation LLM with carefully designed prompts. To improve Self-Instruct, Evol-Instruct Xu et al. (2023) enhances the complexity and diversity of the instruction dataset by prompting ChatGPT with heuristic prompts. OSS-Instruct Wei et al. (2023) queries ChatGPT to generate instruction-output pairs by getting inspiration from real-world code snippets.
Recently, some parameter-efficient fine-tuning techniques have been proposed to use MoE for better instruction tuning. For example, LoRAMoE Dou et al. (2023) and MoCLE Gou et al. (2024) propose MoE-like modules that are constructed with Low-Rank Adaptations (LoRA) to improve instruction tuning, while PESC Wu et al. (2024) proposes to integrate adapters into MoE that are upcycled from dense models. Unlike these works, FT focuses on full-parameter fine-tuning, which is proven generally stronger than parameter-efficient fine-tuning Chen et al. (2022).
@@ 2.3 Weight Averaging
Weight averaging is a commonly used technique to improve the performance of deep learning models. For example, Model Soups Wortsman et al. (2022) averages the weights of multiple models that are initialized from the same pre-trained model but finetuned with different hyperparameter configurations to improve the accuracy and robustness of the fine-tuned model. However, only a few works have been proposed to merge experts of an MoE layer to a normal FFN layer with weight averaging to reduce both parameter and computation overhead of inference. For example, OneS Xue et al. (2022) proposes several simple weight averaging methods to merge expert networks of a BERT-based MoE model. Closely related to our work, Experts Weights Averaging (EWA) Huang et al. (2023) proposes to convert an MoE model to a dense model with two steps: (i) During MoE training, EWA conducts weighted averaging of all the expert weights after each weight update of MoE, which is based on a manually-crafted hyperparameter ; (ii) After training, EWA converts each MoE layer into an FFN layer by uniformly averaging the experts.
Different from all the aforementioned existing works, FT is the first work proposing a learnable mechanism to merge expert networks in the upcycled MoE model. Furthermore, while the training scheme of EWA is deeply coupled to a specific MoE architecture, FT can be easily adapted to different MoE architectures by only adjusting the final merging process. In addition, unlike EWA, FT does not introduce any hyperparameters into the training of large MoE models, significantly reducing the computational resources for hyperparameter searching. Our empirical results in Section 4 also demonstrate the clear advantage of FT.
@@ 3.1 Upcycling
Inspired by sparse upcycling Komatsuzaki et al. (2023), we convert the pre-trained dense LLM to a new MoE by initializing each expert of each MoE layer as a copy of the original FFN layer in the dense model, while directly copying the remaining layers from the dense model to the new MoE model. However, the performance gain brought by sparse upcycling is negligible with a limited training budget Komatsuzaki et al. (2023) – which is exactly the situation we are facing during instruction tuning. Intuitively, it is because each expert in the upcycled MoE model is trained on fewer instruction data than the original dense model does because traditional routers used in sparse upcycling will assign different tokens to different experts and thus reduce the amount of data each expert is trained on Gou et al. (2024). Consequently, inspired by DeepSeekMoE Dai et al. (2024) and MoCLE Gou et al. (2024), FT introduces the shared expert setting into sparse upcycling to tackle this challenge. We further propose a novel routing weight normalization strategy for FT to avoid the potential performance degradation caused by the scale mismatch problem Wu et al. (2022).
@@ 3.1.1 Shared Expert for Upcycling
During upcycling, we isolate one shared expert among all the other normal experts in each MoE layer, where the shared expert will be deterministically assigned to handle all the tokens while other normal experts are assigned by the router. By doing so, the upcycled MoE model can achieve a clear performance boost in instruction tuning, where the shared expert learns general knowledge across the whole instruction dataset while other normal experts learn specific knowledge among different instructions assigned by the router. Formally, the output hidden state of the -th MoE layer when processing the -th token can be expressed as:
where refers to the total number of experts, refers to the number of activated experts, refers to the gate value for the -th expert given the -th token, refers to the -th expert, refers to the output hidden state of the -th attention layer given the -th token (which is also the input of the -th MoE layer), refers to the affinity score between the -th expert and the -th token, refers to the maximum affinity score among all the experts besides the shared expert, refers to a function extracting highest scores out of , refers to a set of highest affinity scores among all the experts besides the shared expert, and refers to the centroid of the -th expert in the -th MoE layer.
@@
Routing Weight Normalization. Although the shared expert setting is also used in recent works Dai et al. (2024); Gou et al. (2024), we cannot directly follow their routing strategy because they cannot handle a scale mismatch problem that is unique for sparse upcycling. The scale mismatch problem is that differences between the scale of the output of the upcycled MoE layer and that of the original FFN layer can cause performance degradation Wu et al. (2022). To handle this problem, we need to ensure the sum of equals 1, so that the output of the MoE layer matches that of the FFN layer in scale. To do so, we normalize the affinity scores of top normal experts with Softmax and scale their sum to to make sure that the sum of the of top experts, including one shared expert and normal experts, equals 1.
@@ 3.2 Merging
We propose a learnable model merging method to convert the large MoE model, namely MoE, back to a dense model FT. By doing so, we expect FT to keep the boosted performance gained during upcycling while keeping its model size the same as the original dense model size to avoid any additional inference overhead. Inspired by Model Soups Wortsman et al. (2022), we choose to merge MoE by learning the mixing coefficients that can be used to average the parameters of all experts in each MoE layer to obtain a normal FFN layer, while directly copying other remaining layers.
Formally speaking, given the weights of experts at the -th layer , the process of merging each MoE layer to an FFN layer can be stated below:
where denotes the merged parameter of all experts and denotes the learnable mixing coefficient of expert . Mixing coefficients is parameterized as the output of a softmax, ensuring that is positive and . Given input , we denote the output of a neural network with parameters as . For loss and instruction dataset , such mixing coefficients of all the layers can be learned via:
where refers to all the remaining layers of MoE other than MoE layers.
While the learning process defined in Eq. (4) is the most intuitive way of learning , our preliminary experiment shows that, due to the shared expert setting, it tends to increase the mixing coefficient of the shared expert at each layer as much as possible to decrease the loss. It is not helpful because, although the shared expert has learned general knowledge across the whole instruction dataset and needs a relatively large mixing coefficient, we still need to keep the scale of the mixing coefficient of other normal experts at a certain level to keep the specific knowledge learned by other normal experts in the merged parameter .
To solve this issue, we introduce a shared expert rate to fix the mixing coefficient of the shared expert and learn the mixing coefficients of the remaining normal experts which sums to in each layer. By doing so, we can easily control the scale of the mixing coefficient of the shared expert, while still being able to learn the optimal layer-wise mixing coefficients of other normal experts. Let’s say is the shared expert of the -th layer, then Eq. (3) and Eq. (4) can be reformulated as follows:
In practice, we uniformly initialize the mixing coefficients of all the normal experts as , which is then trained on the same instruction dataset used during upcycling.